Assume that the short-run cost and demand data given in the tables below confront a monopolistic competitor selling a given product and engaged in a given amount of product promotion.

Cost Data Demand Data
Total Output Total Cost Quantity Demanded Price
0 $25 0 $60
1 40 1 55
2 45 2 50
3 55 3 45
4 70 4 40
5 90 5 35
6 115 6 30

What will the maximum total profits be?

Multiple Choice

  • $90

  • $65

  • $85

  • $110
Answer #1

Answer

MC(n)=TC(n)-TC(n-1)
MC(n)=marginal cost of n th unit
TC(n)=Total cost of n units of output
MC(1)=40-25=15 and so on

TR=P*Q

TR(0)=0*60=0 and so on

Marginal revenue of n th unit of output =total revenue of n unit of output – total revenue of n-1 units of output

MR(n)=TR(n)-TR(n-1)
MR(1)=55-0=55 and so on

Profit=TR-TC

Profit (0)=0-25=-25 and so on

Cost Data Demand Data
Total Output Total Cost Quantity Demanded Price TR MR MC Profit
0 25 0 60 0 -25
1 40 1 55 55 55 15 15
2 45 2 50 100 45 5 55
3 55 3 45 135 35 10 80
4 70 4 40 160 25 15 90
5 90 5 35 175 15 20 85
6 115 6 30 180 5 25 65

the firm maximizes profit at MR=MC or the nearest lower MC

where

Q=4 units, MR=25 and MC=15

Profit=160-70=$90

Option 1